OpenStudy (anonymous):

Evaluate -5x 2 - 3x for x = -2. 14 -14 -26

OpenStudy (dakid88):

C

OpenStudy (andrewherpeche):

c

OpenStudy (retireed):

f(−2)= -5x^2−3x PEDMAS (I learned that here!) There are no parenthesis, so do the exponent first, then the multiplies and finally the subtraction. x = -2 -5(-2^2) - 3(-2) = -5(4) -3(-2) = -20 - (-6) = -20 + 6 = __________

OpenStudy (andrewherpeche):

lets give each other medals

OpenStudy (andrewherpeche):

now u give me one